### v2

@@ Abstract
There is currently no consensus on how the global population will evolve in the next decades and in the next century [1]. The reason for this uncertainty is the absence of reliable population dynamic models. A recent study has shown, by gathering all the global population growth data available since the start of the Holocene to date, that the global population trend vs time in various epochs is mathematically described by four types of different functions [2]. Currently, none of the known mathematical population-dynamic models can reproduce all these four types of functions as its solutions. This is of course a great impediment on the way towards forecasting trends in global population growth, which is an ever more urgent exercise in an era dominated by substantial climate and ecological threats. In this paper, we remedy to this situation by reporting on a population dynamic model, a single nonlinear differential equation adapted from the physics of disordered systems [3], which is able to mathematically describe all the various regimes encountered in the global population recorded as a function of time, over the past 12000 years until now. Regimes of simple exponential growth (Malthus), logistic (Verhulst) plateaus as well as stretched-exponential and compressed-exponential growth regimes are all reliably described by this mathematical equation in its various limits. Besides showing that this is, indeed, the most general population dynamic model, we use it to explore its solutions projected into the future. In particular, two different scenarios are predicted. In one of them, which assumes that the future evolution would continue along a similar pattern as the past decades (hence without any major global ecological crisis affecting the resource exploitation), a von Foerster-type doomsday scenario with a sudden rise of the global population to unsustainable levels could appear as early as 2078. In the opposite scenario, if a global ecological crisis were to set in today, affecting the ability to exploit resources, given the current estimates of the Earth’s carrying capacity, the global population is forecasted to reduce by half by 2064. Furthermore, the proposed dynamic model provides with a new aggregated parameter ($K$, in the model) that can be monitored and controlled so as to avoid the doomsday scenarios.
@@
$y(t)=\frac{1}{K}\textrm{Ei}^{-1}(\textrm{Ei}(1)-t)=\frac{1}{K}\textrm{Ei}^{-1}(1.89512-t)$ (3)
@@
where $\mu=1.4513692...$ is the Ramanujan-Soldner constant and $P_{n}$ are polynomials defined recursively by $P_{0}(x)=x,\ P_{n+1}(x)=x(P_{n}^{\prime}(x)-nP_{n}(x))$.

### v3

@@ Abstract
We show that a simple nonlinear differential equation (originally studied in the physics of disordered systems) is able to mathematically describe the global population growth over the past 12000 years. Different regimes of population growth since the early Neolithic until today are shown to be all solutions to the same nonlinear differential equation in its various limits. These also include the well-known Malthus (exponential) and Verhulst (logistic) growth regimes, as well as von Foerster’s “doomsday” formula. All these limits correspond to neglecting higher-order terms in a more general nonlinear dynamic model described by the proposed nonlinear differential equation. While the older models may provide valid fittings to limited time intervals in the global population growth curve in time, their clearly approximate nature prevents them from being predictive over longer periods of time. The proposed comprehensive solution of the proposed model is instead well suited to provide predictions for future scenarios. These include halving of the global population as early as 2064 due to resource depletion, if the effect of the Earth’s limited carrying capacity were to set in today.
@@
$y(t)=-\frac{1}{K}\textrm{Ei}^{-1}\left(\textrm{Ei}(-K)+t\right)$ (3)
@@
where $\mu=1.4513692...$ is the Ramanujan-Soldner constant and $P_{n}$ are polynomials defined recursively by $P_{0}(x)=x,\ P_{n+1}(x)=x(P_{n}^{\prime}(x)-nP_{n}(x))$. We have checked that the solution produced in this way closely matches the numerical solution (apart from very few discontinuous points which are due to the complexity of numerically inverting the exponential integral function).